A planet hunter in the North Francesco Pepe Geneva Observatory HARPS-N Consortium & the
Consortium and Science Team • Consortium • Geneva Observatory (Head), CfA (Cambridge), Harvard University. INAF-TNG, University of St. Andrews, University of Edinburgh, Queens University Belfast • Many thanks to all the collaborator for the incredible work! • HARPS-N Science Team • Andrew Collier-Cameron, David Charbonneau, David Latham, Mercedes Lopez-Morales, Christophe Lovis, Michel Mayor, Giusi Micela, Francesco Pepe, David Phillips, Giampaolo Piotto, Didier Queloz, Kenb Rice, Dimitar Sasselov, Damien Ségransan, Alessandro Sozzetti, Andrew Szentgyorgyi, Stéphane Udry, Chris A. Watson and Collaborators ...
HARPS-N Historical background • Idea was born in 2005 (Latham, Mayor, Sasselov). The need for a HARPS-like instrument in the northern hemisphere was recognized, to follow-up by RVs the many expected Kepler candidates, only visible from the north ... to be ready in 2009 • PDR in December 2007 • 2008 and 2009 Project stalled because of financial difficulties • Spring 2010 re-organization of the Consortium, the project and the financial resources. • Summer 2010 negotiation with INAF-TNG • December 2010 signature of Agreement with INAF and Collaboration agreement • March 2012 installation at La Palma • August 2012 start of operations • ...
HARPS-N Historical background • September 16th, 2012. Repair of Detector-dewar leak and break-down of CCD • September 30th, 2012. PO for the new CCD ... without money ;-( • November 9th, 2013. New chip installed and operating ... ;-) • February 9th - 18th, 2013. Final mission: • New Sequencer (it is really nice, now;-) • Upgrade of the Labview SW, in particular guiding • Installation of all sensors, isolation, thermal control • Focus procedure using the exposure meter • Improvement of the interfaces with TNG • New Fabry-Pérot vessel: Now stable to better than 0.5m/s per night!!! • Many thanks to all the good guys !!!
HARPS-N Latest Sequencer changes(End of June 2013) • Using look-up table for telescope focus automatically (not the same as through-focus) • Abort during pointing works correctly • Focus procedure improved. It use now the Exposure Meter and the shutter is controlled directly from the Sequencer • Stop tracking after focus • Stop AG after each OB to avoid unwanted telescope tracking • 'Real' lamp and dust cover now updated every 4 seconds • Lamps status checked at the beginning of the OB. • THAx warm-up time of 10 minutes checked at the beginning of each template. • THA1 asked for switch-off after each OB (timer of 5 minutes) • New LED FF template available and executed each night. No DRS yet.
HARPS-N Other improvements • Now overflow from Exposure Meter anymore (since June 27th) • Exposure Meter provides efficiency and mid-time of exposure in fraction instead of time (since April 4th) • Headers complete and uniform (since April 4th). Update of the archive still pending • Problem with ‘axis failed’ solved (since July 2013) • AG stability and reliability improved, especially on faint objects (since may 2013) • FP stable and available for any/faint sources. Just use it by selecting ‘wavesimult’ instead of ‘thosimult’ OBs. DRS will work as usual. Nothing else to do. Recommended for objects fainter than V = 11, if simultaneous reference necessary. Normally stability of the spectrograph better than 1.5 m/s. FP better than 10-20 cm/s per night. • DRS improved. New version not yet released (see later in the presentation) • A few problems (failures) repaired thanks to the excellent actions by Adriano Ghedina and all the TNG friends.
The HARPS-N CCD E2V CCD 231 4k4, 15 µm pixels Graded coating
HARPS-N Still to be done • Update archived files (headers) • Release new DRS (very very soon) • Replace Labview Exposure Meter by Java Exposure Meter (Nicolas Buchschacher) • It will be possible to see the obtained SNR in the Sequencer window. The plan is to have an option to ‘halt on SNR’, in addition to ‘TEXP’ • Replace TH operations by FP operations -> solve the lamp availability problem • Any other idea? Contributions are welcome. • A few issues with telescope reliability.
Efficiency & bench mark mv = 8 G-star in 10 minutes K0V: SNR=60 <=> 1 m/s in 1 hour on mv = 12
RV standard 1.29 m/s rms
Kepler-78b: Radial velocities Pepe et al. 2013 (submitted)
An update on the HARPS-N DRS Christophe Lovis Geneva Observatory HARPS-N Consortium & the
The scientific reduction recipe • Read raw image and ‘format’ frame • Cosmetic correction (interpolation of identified ‘bad’ columns and pixels • BIAS correction using overscan only • Read order profile from spectral FF calibration (profile determined from FF, no analytical function) • Convert to e- • Correct for CTE effects (new) • Dark correction (negligible) • General (low spatial frequencies) background measurement • Optimum extraction (A, B or both) considering background and contamination reference frame (thoB or waveB) • Flat-field correction using FF reference calibration • Apply wavelength calibration lamda(x) determined from ThAr calibration frame on fiber A! • Save e2ds spectra
The scientific reduction recipe • Produce and save s1d spectra (de-instrument-transmittance, de-blaze and join) • Compute and display instrumental drift using fiber B (either ThAr -> ThAr or FP -> FP) • Correct wavelength solution of fiber A from drift measured on fiber B • Compute barycentric velocity and correct wavelength solution • Reject cosmics • Compute flux-weight correction • Compute RV by means of numerical cross correlation CCF (+ BIS, FWHM, Contrast, logR’HK, errors, etc.) • Save CCF files and other products
Latest DRS upgrades • Background and contamination correction implemented and working • CTE correction implemented and working (on new frames only, at present) • Computation of R’HK implemented and tested • Flux-weight correction implemented in RV computation • Correction of mid-time of exposure taken into account since June 2013. • Other adaptation done, e.g. management of changing CCDs
Still to be done • DRS to be released • Update of archive data (headers) and re-run DRS (by archive personnel and TNG) • Many other ideas, but HR needed ... Any contribution will be welcome
Stable IP: BIS, FWHM, CONTRAST KOI-141